40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Feb 16 14:56:44 2021, gautam, Update, SUS, aLIGO Sat Amp installed, powered MC1_deadUL.pngIMG_9100.jpg
    Reply  Wed Feb 17 16:11:53 2021, gautam, Update, SUS, aLIGO Sat Amp installed, powered 
       Reply  Thu Feb 18 15:33:21 2021, gautam, Update, SUS, aLIGO Sat Amp installed, powered and commissioned 
          Reply  Thu Feb 18 18:05:04 2021, gautam, Update, SUS, aLIGO Sat Amp characterization OSEMnoise.pdfLEDdriveNoise.pdf
             Reply  Thu Feb 18 20:20:25 2021, Koji, Update, SUS, aLIGO Sat Amp characterization 
Message ID: 15814     Entry time: Wed Feb 17 16:11:53 2021     In reply to: 15809     Reply to this: 15817
Author: gautam 
Type: Update 
Category: SUS 
Subject: aLIGO Sat Amp installed, powered 

There is some non-trivial sign flipping in the sensors/coils in this new setup because it is a hybrid one with the old interfacing electronics (D000210, D010001) and the new Satellite Amplifier (D080276). So I haven't yet gotten the damping working. I am leaving the PSL shutter closed and will keep working on this today/tomorrow. I have made various changes to the c1mcs realtime model and the c1susaux database record where MC1 is concerned. I have backups of the old ones so we can always go back to that if we so desire.

In the meantime, the PSL shutter is closed and there is no light to the IFO.


Update 1700: I've implemented some basic damping and now the IMC is now locked. The WFS loop runs away when I enable it, probably some kind of weird interaction with the (as of now untuned) MC1 local damping loops. I will write up a more detailed report later.


Update 2300: Did the following:

  1. Re-calibrated the cts2um filter in all SENSOR filter banks to account for the increased transimpedance and LED drive current. I judged the overall scaling to be x0.25 but this can be calibrated against the bounce peak height for example (it lines up pretty well).
  2. Re-measured the input matrix - it was very different from what was loaded. I am measuring this again overnight for some consistency.
  3. Re-tuned the damping gains. Now the optic damps well, and the loops seem file to me, both via broadband noise injection TF and by step response metrics.
  4. Yet, the WFS servo cannot be enabled. The WFS signal is summed in before the output matrix so I don't know why this would have a different behavior compared to the local damping, or indeed, why this has to be changed. Will need some (WFS) sensing/actuation matrix measurements to know more.

Dropping this for tonight, I'll continue tomorrow. Meanwhile, the OSEM input matrix measurement is being repeated overnight. PSL shutter is closed.

ELOG V3.1.3-